OpenStudy (anonymous):

Find the limit. Lim (sin2x) x->0 (sin3x)

OpenStudy (anonymous):

trig id: \[\sin(2x) = 2\sin(x)\cos(x)\] trig id: \[\sin(3x) = -4(\sin(x))^3 + 3\sin(x)\] you will to some reduction then apply the limits. this should put you in good start.

OpenStudy (anonymous):

Ok, thanks a lot for the help. :)